Operations Report – February 1st

Crew 188 Operational Report 01/02/2018
Sol 4
Zac Trolley
Non-Nominal Systems: Propane supply, Microwave, Rover Charging infrastructure, Greenhab temperature sensor. Notes on non-nominal systems: See individual systems at the end of the report. Generator (hours run): 18.8h
Solar: SOC 81%
Diesel: 78%
Propane: 62 psi
Ethanol Free Gasoline (5 Gallon containers for ATV): 8 Gallons Water (trailer): 0 Gallons
Water (static): 300 Gallons
Trailer to Static Pump used: No
Water (loft) – Static to Loft Pump used: Yes
Meter 130607.8
Toilet tank emptied: Yes

ATVs Used: None
Oil Added: None
ATV Fuel Used: None
# Hours the ATVs were Used today: 0
Notes on ATVs: none
Deimos rover used: No
Hours: Hours: 106.9h (Accurate as of two days ago, rover was not on campus during EVA to check hours.)

Beginning charge: N/A
Ending charge: N/A
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: no
Hours: 17.5h
Beginning charge: N/A
Ending charge: N/A
Currently charging: No

Opportunity rover used: no
Hours: 13.5
Beginning charge: N/A
Ending charge: N/A
Currently charging: No

Curiosity rover used: No
Hours: 9.3h
Beginning charge: N/A
Ending charge: N/A
Currently charging: No

HabCar: Not used
General notes and comments: The Hab Director mentioned that the Rovers required an extension cord that is shorter than 25 feet. We estimate the cords being used now are 50 foot cables. I was able to find a 25 foot extension cable, meaning we could charge one rover at time based on the above specifications. Can we have the appropriate extension cables sent to the Hab in order to charge the Rovers?
Also, the configuration the rovers are in means that the rovers need to back out before they are used. Backing up in an EVA suit is hazardous since you can’t see behind you in a suit. Please advise if there is another configuration that allows charging with a 25 foot extension cable as well as the ability to drive forward during EVAs.

Summary of internet: Internet connectivity continues to be a problem even with the data purchased. Devices are having troubles connecting and
Summary of suits and radios: There is a bent clasp on suit #1 on the helmet collar. It was probably bent during the EVA, the system is still functional.
Summary of Hab operations: The microwave isn’t working anymore. The buttons work, the time counts down, but the microwave doesn’t actually produce waves to heat anything.
Summary of EVA operations: We are still brainstorming ways to combat the helmet fog. Without proper ventilation under the helmet ring, the facemasks rapidly fog up. It is becoming a safety issue on EVA.
Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.
Summary of Science Dome operations: The science dome is awaiting a portable heater as per the Hab Director. Summary of RAM operations: Not Operational.
Summary of health and safety issues: The propane detector has arrived, and there is a positive detection of combustible gas in the hab.
We calibrated the sensor outside, and moving the detector in different places in the Hab clearly shows that there are areas where there is a higher concentration of combustible gas than others. The instrument is of low quality and does not show us the ppm, so we have no way to tell what the levels are, and if they are dangerous.
The leak was found using soapy water, and the fitting was tightened by the crew. We will investigate tomorrow and hopefully the leak is fixed. Questions, concerns and requests to Mission Support: All nominal [end]

Copyright © The Mars Society. All rights reserved. | Main Site